What is the equation of a line perpendicular to 2x-8y=3 that contains the point (-8, 11)?
Answer by Alan3354(69443)

What is the equation of a line perpendicular to 2x-8y=3 that contains the point (-8, 11)?
-----------------------
Step 1, find the slope of the given line, m1.
The slope of lines perpendicular is the negative inverse of m1.
Call it m.
-------
Step 2, use y-y1 = m*(x-x1) where (x1,y1) is the point.
----
I don't need the practice, you do.
-----
PS Not "the equation," an equation.